A weather balloon with a volume of 170.0 L is launched at 20 degrees C at sea level, where the atmospheric pressure is 1.00 atm. The balloon rises to an altitude of 20,000 m, Where atmospheric pressure is 59 mmHg and the temperature is 210 K. What is the volume of the balloon at 20,000 m?
Answer #1

P1 = 1atm = 760mmhg

V1 = 170L

T1 = 20C = 20+273 = 293K

P2 = 59mmhg

V2 =

T2 = 210K

P1V1/T1 = P2V2/T2

V2 = P1V1T2/P2T1 = 760*170*210/59*293 = 1569.5L